OpenStudy (anonymous):

write the equation of the line parallel to y=-3, passing through the point (2,5)

OpenStudy (anonymous):

y=(0/1)2+5

OpenStudy (anonymous):

leme check...im not sure

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

well, since the first line is just y = 3...that means the line is going horizontal so excuse me its (1,0)

OpenStudy (anonymous):

m is the slope...equation is y+mx+b

OpenStudy (anonymous):

*y=mx+b

OpenStudy (anonymous):

this doesnt have a slope just the y intercept. So what would be the final answer im confused

OpenStudy (anonymous):

5=(1/0)2+b

OpenStudy (anonymous):

...wait a sec... leme check my books

OpenStudy (anonymous):

its been years since i have done this

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

this will take 5 minutes at the most

OpenStudy (anonymous):

thats fine

OpenStudy (whpalmer4):

\[y = -3 \]is a horizontal line. It therefore has \(m = 0\) \[y = mx + b\]\[y = mx - 3\]\[y = 0x-3\]\[y = -3\]

OpenStudy (anonymous):

...he will answer it...its 10 oclock where i am an i am tierd...sorry i held u up

OpenStudy (whpalmer4):

To fit a line through the desired point with the desired slope, use the point-slope formula: \[y-y_0 = m(x-x_0)\] Your target point is \((x_0,y_0)\), and your slope is \(m = 0\)

OpenStudy (anonymous):

oh I see Thanks to both of you! :)

OpenStudy (whpalmer4):

The final result here is rather trivial, of course: \[y - 5 = 0(x-2)\]\[y -5 = 0\]\[y = 5\]